# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

### v2

@@ Abstract
Recent text-to-image models generate high-quality images from text prompts but lack precise control over specific components within visual concepts. Therefore, we introduce component-controllable personalization, a new task that allows users to customize and reconfigure individual components within concepts. This task faces two challenges: semantic pollution, where undesirable elements distort the concept, and semantic imbalance, which leads to disproportionate learning of the target concept and component. To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics. The experimental results show that MagicTailor outperforms existing methods in this task and enables more personalized, nuanced, and creative image generation.
Figure 1: (a) Personalization Illustration: Showing how text-to-image diffusion models learn and generate a visual concept from reference images. (b) Component-Controllable Personalization Illustration: Introducing a task that aims to modify specific components of visual concepts during personalization. (c) Generated Images by MagicTailor: Demonstrating MagicTailor’s effectiveness in component-controllable personalization, with red and blue circles highlighting the target concept and component, respectively.
@@ 1 Introduction
Text-to-image (T2I) diffusion models [35, 39, 31, 5] have shown impressive capabilities in generating high-quality images from textual descriptions. While these models can generate images that align well with provided prompts, they struggle when certain visual concepts are hard to express in natural language. To address this, methods like [11, 36] enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images. This process, as shown in Fig. 1 (a), is referred as personalization.
However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control over their components. This limitation hinders their practical use in real-world applications. Since concepts often consist of multiple components, such as a house with walls, windows, and doors, a key problem in personalization is how to effectively control and manipulate these individual components.
Figure 2: Major challenges in component-controllable personalization. (a) Semantic pollution: (i) Undesired elements may interfere with the personalized concept. (ii) A simple mask-out strategy causes unintended results, while (iii) DM-Deg effectively suppresses unwanted semantics. (b) Semantic imbalance: (i) Simultaneously learning the concept and component can distort either one. (ii) DS-Bal ensures balanced learning, improving personalization. (c) Identity fidelity performance: Calculating DreamSim [10] scores on a collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
@@
One key challenge of this task is semantic pollution (Fig. 2 (a)), where unwanted visual elements inadvertently appear in the generated image, “polluting” the personalized concept. This happens because the T2I model often mixes visual semantics from different regions during training. Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and creates unintended compositions. Another challenge is semantic imbalance (Fig. 2 (b)), where the model overemphasizes certain aspects, leading to unfaithful personalization. This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
@@
To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models ( Fig. 1 (c)). We first use a text-guided image segmenter to generate segmentation masks for both the concept and component and then design Dynamic Masked Degradation (DM-Deg) to transform reference images into randomly degraded versions, perturbing undesired visual semantics. This method helps suppress the model’s sensitivity to irrelevant details while preserving the overall visual context, effectively mitigating semantic pollution. Next, we initiate a warm-up phase for the T2I model, training it on the degraded images using a masked diffusion loss to focus on the desired semantics and a cross-attention loss to strengthen the correlation between these semantics and pseudo-words. To address semantic imbalance, we develop Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm that balances the learning of visual semantics. In this phase, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preservation regularization. This ensures more accurate and faithful personalization of the target concept and component.
We validate the effectiveness of MagicTailor through various qualitative and quantitative experiments, demonstrating state-of-the-art (SOTA) performance in component-controllable personalization. Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor. Additionally, we show its potential for enabling a wide range of creative applications.
@@ 2 Related Works
Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions. Early methods employed Generative Adversarial Networks (GANs) [33, 53, 28, 59], and transformers [6, 30, 7, 55, 4] also showed the potential in conditional generation. The advent of diffusion models has ushered in a new era in T2I generation [20, 39, 31, 5, 54]. Leveraging these models, a range of related applications has rapidly emerged, including image editing [20, 24, 16, 9, 3], image completion and translation [50, 49, 22], and controllable image generation [56, 45, 58]. Despite advancements in T2I diffusion models, generating images that accurately reflect specific, user-defined concepts remains a challenge. This study explores component-controllable personalization, which allows precise adjustment of specific concepts’ components using visual references.
@@
Personalization seeks to adapt T2I models to generate given concepts using reference images. Initial approaches [11, 36] addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model. Additionally, low-rank adaptation (LoRA) [15] has been widely adopted by the research community for personalization [37], offering an efficient and lightweight solution. The scope of personalization has expanded to encompass multiple concepts [19, 2, 13, 12, 25]. Besides, several studies [48, 21, 40, 46, 44] have explored tuning-free approaches for personalization, but these often necessitate training an encoder on large, domain-specific image datasets [57]. In this work, MagicTailor adopts the widely used paradigm of test-time optimization, providing stable performance and precise control over personalization.
@@ 3 Methodology
Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$. In this work, we focus on a practical setting involving one concept and one component. Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear). In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers. The goal of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and the component from $\mathcal{I}$. Using text prompts with $\mathcal{P}$, the fine-tuned model should generate images that integrate the personalized concept with the specified component.
This section begins by providing an overview of the MagicTailor pipeline in Sec. 3.1 and then delves into its two core techniques: Dynamic Masked Degradation in Sec. 3.2 and Dual-Stream Balancing in Sec. 3.3.
@@ 3.1 Overall Pipeline
Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity can mitigate noise memorization. (c) We report IQA results of Q-Align [47] on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
@@
where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$. For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) [15] manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen. Thereafter, we employ Dual-Stream Balancing (DS-Bal) to address semantic imbalance. In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other samples.
@@ 3.2 Dynamic Masked Degradation
Semantic pollution is a major challenge for component-controllable personalization. As shown in Fig. 2 (a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person. Masking regions outside the target concept and component can damage the overall context, leading to overfitting and odd compositions (Fig. 2 (a.ii)). To address this, undesired visual semantics in reference images must be handled appropriately. We propose Dynamic Masked Degradation (DM-Deg), which dynamically perturbs undesired semantics to suppress their influence on the T2I model while preserving the overall visual context (Fig. 2 (a.iii)&(c)).
@@
where $\odot$ denotes element-wise multiplication, and $\alpha_{d}\in[0,1]$ is a dynamic weight controlling the degradation intensity. While previous works [48, 21] have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context. By introducing $\hat{I}_{nk}$, we can suppress the T2I model from perceiving undesired visual semantics in out-of-mask regions, as these semantics are perturbed by random noise at each training step.
@@
Unfortunately, the T2I model may gradually memorize the introduced noise while learning meaningful visual semantics, leading to noise appearing in generated images (Fig. 4 (a)). This behavior is consistent with previous observations on deep networks [1]. To address this, we propose a descending scheme that dynamically regulates the intensity of the imposed noise during training. This scheme follows an exponential curve, maintaining a relatively high intensity in the early stages and decreasing it sharply in later stages. Let $d$ denote the current training step and $D$ denote the total training step. The curve of dynamic intensity is defined as:
@@
where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ controls the descent rate. We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$, tuned within the powers of $2$. This dynamic intensity scheme effectively prevents semantic pollution and mitigates the memorization of introduced noise, leading to improved generation performance (Fig. 4 (b)&(c)).
@@ 3.3 Dual-Stream Balancing
Another challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component. Concepts generally have richer visual semantics than components (e.g. person vs. hair), but in some cases, the component may have more complex semantics (e.g., simple tower vs. intricate roof). This imbalance complicates joint learning, leading to overemphasis on either the concept or the component, and resulting in incoherent generation (Fig. 5 (a)). To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm with online and momentum denoising U-Nets (Fig. 3) for balanced semantic learning, improving personalization fidelity (Fig. 5 (b) & Fig. 2 (c)).
@@
where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization [32]. The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component. Such an optimization scheme can effectively modulate the learning dynamics of different samples and avoid the overemphasis on any particular one.
@@
At a training step, the sample neglected in $\mathcal{L}_{\text{diff-max}}$ may suffer from knowledge forgetting. This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others. In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net to preserve the learned visual semantics of the other sample in each training step. Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which is expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set. Then, we use the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
@@
where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA [42] with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step. By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other samples while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$. Overall, DS-Bal can be considered a mechanism to adaptively assign target labels $\epsilon_{n}$ or preserving labels $\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ to different samples, enabling dynamic loss supervision (Fig. 3). Using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as:

### v3

@@ Abstract
Text-to-image diffusion models can generate high-quality images but lack fine-grained control of visual concepts, limiting their creativity. Thus, we introduce component-controllable personalization, a new task that enables users to customize and reconfigure individual components within concepts. This task faces two challenges: semantic pollution, where undesired elements disrupt the target concept, and semantic imbalance, which causes disproportionate learning of the target concept and component. To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics. The experimental results show that MagicTailor achieves superior performance in this task and enables more personalized and creative image generation.
@@ 1 Introduction
Text-to-image (T2I) diffusion models (Rombach et al., 2022; Ramesh et al., 2022; Chen et al., 2023) have shown impressive capabilities in generating high-quality images from textual descriptions. While these models can generate images that align well with provided prompts, they struggle when certain visual concepts are hard to express in natural language. To address this, methods like (Gal et al., 2022; Ruiz et al., 2023) enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images. This process, as shown in Fig. 1(a), is referred as personalization.
However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control of these concepts. Such a limitation hinders their practical use in real-world applications, restricting their potential for creative expression. Inspired by the observation that concepts often comprise multiple components, a key problem in personalization lies in how to effectively control and manipulate these individual components.
@@
Figure 1: (a) Personalization: T2I models learn from reference images and then generate predefined visual concepts. (b) Component-controllable personalization: T2I models learn from additional visual references and then enable the integration of specific components into given concepts, further unleashing creativity. (c) Generated images by MagicTailor: MagicTailor can effectively achieve component-controllable personalization. Note that red and blue circles indicate the target concept and component, respectively.
One challenge of this task is semantic pollution (Fig. 2(a)), where unwanted visual elements inadvertently appear in generated images, “polluting” the personalized concept. This happens because the T2I model often mixes visual semantics from different regions during training. Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and causes unintended compositions. Another challenge is semantic imbalance (Fig. 2(b)), where the model overemphasizes certain aspects, leading to unfaithful personalization. This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
Figure 2: Major challenges in component-controllable personalization. (a) Semantic pollution: (i) Undesired elements may interfere with the personalized concept. (ii) A simple mask-out strategy causes unintended results, while (iii) DM-Deg effectively suppresses unwanted semantics. (b) Semantic imbalance: (i) Simultaneously learning the concept and component can distort either one. (ii) DS-Bal ensures balanced learning, improving personalization. (c) Identity fidelity performance: Calculating DreamSim (Fu et al., 2023) scores on our collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
@@
To address these challenges, we propose MagicTailor, a novel framework that enables component-controllable personalization for T2I models (Fig. 1(c)). We first use a text-guided image segmenter to generate segmentation masks for both the concept and component and then design Dynamic Masked Degradation (DM-Deg) to transform reference images into randomly degraded versions, perturbing undesired visual semantics. This method helps suppress the model’s sensitivity to irrelevant details while preserving the overall visual context, effectively mitigating semantic pollution. Next, we initiate a warm-up phase for the T2I model, training it on the degraded images using a masked diffusion loss to focus on the desired semantics and a cross-attention loss to strengthen the correlation between these semantics and pseudo-words. To address semantic imbalance, we develop Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm that balances the learning of visual semantics. In this phase, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preservation regularization. This ensures more faithful personalization of the target concept and component, resulting in outputs that better align with the intended objective.
In the experiments, we validate the superiority of MagicTailor through various qualitative and quantitative comparisons, demonstrating its state-of-the-art (SOTA) performance in component-controllable personalization. Moreover, detailed ablation studies and analysis further confirm the effectiveness of MagicTailor. In addition, we also show its potential for enabling a wide range of creative applications.
@@ 2 Related Works
T2I generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions. Early methods employed Generative Adversarial Networks (GANs) (Reed et al., 2016; Xu et al., 2018), and transformers (Ding et al., 2021; Yu et al., 2022; Bai et al., 2024) also showed the potential in conditional generation. The advent of diffusion models has ushered in a new era in T2I generation (Li et al., 2024; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023; Xue et al., 2024). Leveraging these models, a range of related applications has rapidly emerged, including image editing (Li et al., 2024; Mou et al., 2024; Huang et al., 2024b; Feng et al., 2024; Huang et al., 2024a), image completion and translation (Xie et al., 2023b, a; Lin et al., 2024), and controllable image generation (Zhang et al., 2023; Wang et al., 2024; Zheng et al., 2023). Despite advancements in T2I diffusion models, generating images that accurately reflect specific, user-defined concepts remains a challenge. This study explores component-controllable personalization, which allows precise adjustment of specific concepts’ components using visual references.
@@
Personalization seeks to adapt T2I models to generate given concepts using reference images. Initial approaches (Gal et al., 2022; Ruiz et al., 2023) addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model. Additionally, low-rank adaptation (LoRA) (Hu et al., 2021) has been widely adopted in this field (Ryu, 2022), providing an efficient solution. The scope of personalization has expanded to encompass multiple concepts (Kumari et al., 2023; Avrahami et al., 2023; Gu et al., 2024; Ng et al., 2025). Besides, several studies (Li et al., 2023; Wei et al., 2023; Zhang et al., 2024b; Song et al., 2024) have explored tuning-free approaches for personalization, but these necessitate additional training on large-scale image datasets (Zhang et al., 2024a). In contrast, MagicTailor is a tuning-base method that requires only a few images and leverages test-time optimization to enable stable performance. Notably, several works (Huang et al., 2024c; Safaee et al., 2024; Ng et al., 2025) have also explored how to learn and customize fine-grained elements. However, these methods can only combine elements or process one element at the same level. By comparison, MagicTailor is a versatile framework able to handle both component-level and concept-level elements.
@@ 3 Methodology
Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ denote a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$. In this work, we focus on a practical setting involving one concept and one component. Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear). In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers. The goal of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and the component from $\mathcal{I}$. Using text prompts with $\mathcal{P}$, the fine-tuned model should generate images that integrate the personalized concept with the specified component.
This section begins by providing an overview of the MagicTailor pipeline in Sec. 3.1 and then delves into its two core techniques in Sec. 3.2 and Sec. 3.3.
@@ 3.1 Overall Pipeline
where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$. For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) (Hu et al., 2021) manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen. Thereafter, we employ Dual-Stream Balancing (DS-Bal) to address semantic imbalance. In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other samples.
@@ 3.2 Dynamic Masked Degradation
Semantic pollution is a significant challenge for component-controllable personalization. As shown in Fig. 2(a.i), the target concept (i.e., person) can be distorted by the owner of the target component (i.e., eye), resulting in a hybrid person. Masking regions outside the target concept and component can damage the overall context, leading to overfitting and odd compositions (Fig. 2(a.ii)). To address this, undesired visual semantics in reference images must be handled appropriately. We propose Dynamic Masked Degradation (DM-Deg), which dynamically perturbs undesired semantics to suppress their influence on the T2I model while preserving the overall visual context (Fig. 2(a.iii)&(c)).
Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity can mitigate noise memorization. (c) We report IQA results of Q-Align (Wu et al., 2023) on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
@@
where $\odot$ denotes element-wise multiplication, and $\alpha_{d}\in[0,1]$ is a dynamic weight controlling the degradation intensity. While previous works (Xiao et al., 2023; Li et al., 2023) have used noise to fully cover the background or enhance data diversity, DM-Deg aims to produce a degraded image $\hat{I}_{nk}$ that retains the original visual context. By introducing $\hat{I}_{nk}$, we can suppress the T2I model from perceiving undesired visual semantics in out-of-mask regions, as these semantics are perturbed by random noise at each training step.
@@
Unfortunately, the T2I model may gradually memorize the introduced noise while learning meaningful visual semantics, leading to noise appearing in generated images (Fig. 4(a)). This behavior is consistent with previous observations on deep networks (Arpit et al., 2017). To address this, we propose a descending scheme that dynamically regulates the intensity of the imposed noise during training. This scheme follows an exponential curve, maintaining a relatively high intensity in the early stages and decreasing it sharply in later stages. Let $d$ denote the current training step and $D$ denote the total training step. The curve of dynamic intensity is defined as:
@@
where $\alpha_{\text{init}}$ is the initial value of $\alpha_{d}$ and $\gamma$ controls the descent rate. We empirically set $\alpha_{\text{init}}=0.5$ and $\gamma=32$, tuned within the powers of $2$. This dynamic intensity scheme effectively prevents semantic pollution and significantly mitigates the memorization of introduced noise, leading to improved generation performance (Fig. 4(b)&(c)).
@@ 3.3 Dual-Stream Balancing
Another key challenge is semantic imbalance, which arises from the disparity in visual semantics between the target concept and its component. Specifically, concepts generally possess richer visual semantics than components (e.g., person vs. hair), but in some cases, components may have more complex semantics (e.g., simple tower vs. intricate roof). This imbalance complicates joint learning, leading to overemphasis on either the concept or the component, and resulting in incoherent generation (Fig. 5(a)). To address this, we design Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm integrated with online and momentum denoising U-Nets (Fig. 3) for balanced semantic learning, aiming to improve personalization fidelity (Fig. 5(b) & Fig. 2(c)).
@@
where minimizing $\mathcal{L}_{\text{diff-max}}$ can be considered as a form of min-max optimization (Razaviyayn et al., 2020). The learning objective of $\epsilon_{\theta}$ may switch across different training steps and is not consistently dominated by the concept or component. Such an optimization scheme can effectively modulate the learning dynamics of multiple samples and avoid the overemphasis on any particular one.
Figure 6: Qualitative comparisons. We present images generated by MagicTailor and other methods across various domains. MagicTailor achieves better text alignment, identity fidelity, and generation quality. Due to space limitations, please zoom in for a better view. More results are provided in Appendix D.
@@
At a training step, the sample neglected in $\mathcal{L}_{\text{diff-max}}$ may suffer from knowledge forgetting. This is because the optimization of $\mathcal{L}_{\text{diff-max}}$, which aims to enhance the knowledge of a specific sample, could inadvertently overshadow the knowledge of the others. In light of this, DS-Bal meanwhile exploits the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ to preserve the learned visual semantics of the other sample in each training step. Specifically, we first select the sample that is excluded in $\mathcal{L}_{\text{diff-max}}$, which is expressed as $S=\{n|n=1,...,N\}-\{n_{\text{max}}\},$ where $n_{\text{max}}$ is the index of the target sample in $\mathcal{L}_{\text{diff-max}}$ and $S$ is the selected index set. Then, we use $\tilde{\epsilon}_{\theta}$ to apply regularization for $S$, with the masked preserving loss as:
@@
where $\tilde{\epsilon}_{\theta}$ is updated from $\epsilon_{\theta}$ using EMA (Tarvainen and Valpola, 2017) with the smoothing coefficient $\beta=0.99$, thereby sustaining the prior accumulated knowledge of $\epsilon_{\theta}$ in each training step. By encouraging the consistency between the output of $\epsilon_{\theta}$ and $\tilde{\epsilon}_{\theta}$ in $\mathcal{L}_{\text{pres}}$, we can facilitate the knowledge maintenance of the other samples while learning a specific sample in $\mathcal{L}_{\text{diff-max}}$. Overall, DS-Bal can be considered a mechanism to adaptively assign target labels $\epsilon_{n}$ or preserving labels $\tilde{\epsilon}_{\theta}(z_{nk}^{(t)},t,e_{n})$ to different samples, enabling dynamic loss supervision (Fig. 3). Using a loss weight $\lambda_{\text{pres}}=0.2$, the total loss of the DS-Bal stage is formulated as: